OpenStudy (anonymous):

Medal? Jess uses her office scanner to scan pages at the rate of $0.09 per page. She decides to rent a scanner for $70 a year. The cost of scanning using the rented scanner is $0.02 per page. Part A: Write an inequality that can be used to calculate the number of pages that Jess should scan in a year so that the amount she pays for the rented scanner is less than the office scanner. Define the variable used. Part B: How many pages should Jess scan in a year to justify renting the scanner? Show your work.

OpenStudy (aripotta):

Part A: 70 + 0.02x < 0.09x can you solve for x and do part b by yourself?

OpenStudy (anonymous):

O = 0.09x R = 70 + 0.02x O is office cost, R is rented cost and x is the number of pages. to find the number of pages that justify, make O = R and solve for x. 0.09x = 70 + 0.02x x = 1000

OpenStudy (anonymous):

need anything else?

OpenStudy (anonymous):

Aballer I'll give you a fan and Ari a medal. And Ari, part b is what confuses me the most.

OpenStudy (aripotta):

70 + 0.02x < 0.09x subtract 0.02x from both sides 70 < 0.07x divide both sides by 0.07 1000 < x so x would have to be over 1000 to justify renting it

OpenStudy (anonymous):

Ohhhh I see now. Thank you for the help, Ari and baller!

OpenStudy (aripotta):

no problem :)